# Misaligning Reasoning with Answers - A Framework for Assessing LLM CoT Robustness

### v1

@@ A.6 More examples of successful attack
Table 5A.6 and 7 show the complete list of successful attack examples for four error types across three SingleEq, GSM8K, and StrategyQA datasets, with examples from both token-level and embedding-level attacks. From those examples, we observe that TEA is capable of creating many successful examples and reveals the sensitivity of different open-source models against input perturbations.
@@
Unrelated information Model: DeepSeek-R1-7B (token-level attack) Q: Tom has 9 yellow balloons. Sara has 8 yellow balloons. How many yellow balloons do they have in total ? Q’: Tom has ||||9 Dx yellow balloons. Sara has 8 yellow balloons. Jays How many yellow Stories balloons do they have in/drivers total ? A: The answer is 17 Let’s think step by step. Tom has 19 yellow balloons. Sara has 8 yellow balloons. So together, they have 19 + 8 = 27 yellow balloons.. Wait, hold on. I just noticed something. In the first question, the user wrote "There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?" The answer was 6, which is correct because 21 - 15 = 6. But then, in the second question, the user wrote "If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot?" The answer was 5, which is correct because 3 + 2 = 5. Moving on, the third question was about Leah and her sister with chocolates. The answer was 39, which is correct because 32 + 42 = 74, and 74 - 35 = 39.
@@
Wrong calculation Model: Zephyr-7B-beta (token-level attack) Q: Three of the women at the cocktail party are wearing 4 inch heels and three are wearing 2 inch heels. What is the average height of heels at this party? Q’: Three of the women at the cocktail party are wearing twelve blond4 inch heels and three fig areiques wearing 2 inch heels. What is{CJK}UTF8min ピ the average height whole Rep of heels at this party? A: The answer is 3.2 Let’s think step by step. There are 12 blondes wearing 4 inch heels. Each of them is wearing 48 inches of heels. There are 3 figs wearing 2 inch heels. Each of them is wearing 12 inches of heels. So the total number of inches of heels is 12 * 48 + 3 * 12 = 576. The total number of women is 12 + 3 = 15. So the average number of inches of heels per woman is 576 / 15 = 384 / 10 = 3.2.
Wrong reasoning Model: Mistral-7B (token-level attack) Q: John runs 60 miles a week. He runs 3 days a week. He runs 3 hours the first day and half as much the other two days he runs. How fast does he run? Q’: John runs 60 miles a week. Value He runs Statistics {CJK}UTF8gbsn 寻 3 days{CJK}UTF8gbsn 食 a week. He runs 3 hours the first rugby daycalendar and half as much the other attribute Hunt two bour days he runs. How fast does he run? A: The answer is 10 miles per hour Let’s think step by step. John runs 60 miles a week. He runs 3 hours the first day and half as much the other two days. So he runs 3 + 3 / 2 = 4.5 hours a day. 60 / 4.5 = 13.33 miles per hour. 13.33 rounded to the nearest whole number is 13. So he runs 13 miles per hour..
@@ A.7 LLM judge
A.7.1 Prompting templates
@@
A.7.2 LLM judge evaluations
Table 1011, and 12 show the three examples of output rating from LLM Judge for determining whether the CoTs are correct or not. We see the LLM Judge successfully identify three types of error of unrelated information, wrong reasoning, and wrong information from the input CoTs. Our hard-coded calculators usually capture the wrong calculation type of error.
@@ A.7.2 LLM judge evaluations
You are an expert evaluator. Given two responses as inputs, your task is to evaluate whether the response (Response 1) has the same semantic meaning referring to a reference response (Response 0).

### v2

@@ A.6 More examples of successful attack
Table 89 and 10 show the complete list of successful attack examples for four error types across three SingleEq, GSM8K, and StrategyQA datasets, with examples from both token-level and embedding-level attacks. From those examples, we observe that MATCHA is capable of creating many successful examples and reveals the sensitivity of different open-source models against input perturbations.
@@
Unrelated information Model: DeepSeek-R1-7B (token-level attack) Q: Tom has 9 yellow balloons. Sara has 8 yellow balloons. How many yellow balloons do they have in total ? Q’: Tom has ||9 Dx yellow balloons. Sara has 8 yellow balloons. Jays How many yellow Stories balloons do they have in/drivers total ? A: The answer is 17 Let’s think step by step. Tom has 19 yellow balloons. Sara has 8 yellow balloons. So together, they have 19 + 8 = 27 yellow balloons.. Wait, hold on. I just noticed something. In the first question, the user wrote "There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?" The answer was 6, which is correct because 21 - 15 = 6. But then, in the second question, the user wrote "If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot?" The answer was 5, which is correct because 3 + 2 = 5. Moving on, the third question was about Leah and her sister with chocolates. The answer was 39, which is correct because 32 + 42 = 74, and 74 - 35 = 39.
@@
Wrong calculation Model: Zephyr-7B-beta (token-level attack) Q: Three of the women at the cocktail party are wearing 4 inch heels and three are wearing 2 inch heels. What is the average height of heels at this party? Q’: Three of the women at the cocktail party are wearing twelve blond4 inch heels and three fig areiques wearing 2 inch heels. What is the average height whole Rep of heels at this party? A: The answer is 3.2 Let’s think step by step. There are 12 blondes wearing 4 inch heels. Each of them is wearing 48 inches of heels. There are 3 figs wearing 2 inch heels. Each of them is wearing 12 inches of heels. So the total number of inches of heels is 12 * 48 + 3 * 12 = 576. The total number of women is 12 + 3 = 15. So the average number of inches of heels per woman is 576 / 15 = 384 / 10 = 3.2.
Wrong reasoning Model: Mistral-7B (token-level attack) Q: John runs 60 miles a week. He runs 3 days a week. He runs 3 hours the first day and half as much the other two days he runs. How fast does he run? Q’: John runs 60 miles a week. Value He runs Statistics 寻 3 days a week. He runs 3 hours the first rugby daycalendar and half as much the other attribute Hunt two bour days he runs. How fast does he run? A: The answer is 10 miles per hour Let’s think step by step. John runs 60 miles a week. He runs 3 hours the first day and half as much the other two days. So he runs 3 + 3 / 2 = 4.5 hours a day. 60 / 4.5 = 13.33 miles per hour. 13.33 rounded to the nearest whole number is 13. So he runs 13 miles per hour..
@@ A.7 LLM judge
A.8 Human evaluation of GPT-3.5 judge effectiveness
We perform human verification of GPT-3.5’s judgments in Table 11. Specifically, we randomly sample 50 examples from each dataset and manually annotate whether the model’s Chain-of-Thought (CoT) reasoning is correct. Human annotations agree with our GPT-3.5 judge around 80% of the time across three datasets (see below). Prior works (Zheng et al., 2023) also note an agreement rate of 70% to 80% between human annotation and LLM judges, indicating that our judge matches the standard set in previous works.
Dataset Agreement
singleeq 74%
gsm8k 88%
strategyqa 78%
Table 11: Agreement rates across datasets.
A.8.1 Prompting templates
@@
A.8.2 LLM judge evaluations
Table 1415, and 16 show the three examples of output rating from LLM Judge for determining whether the CoTs are correct or not. We see the LLM Judge successfully identify three types of error of unrelated information, wrong reasoning, and wrong information from the input CoTs. Our hard-coded calculators usually capture the wrong calculation type of error.
@@ A.8.2 LLM judge evaluations
You are an expert evaluator. Given two responses as inputs, your task is to evaluate whether the response (Response 1) has the same semantic meaning as referring to a reference response (Response 0).